# Meta-DAN: towards an efficient prediction strategy for page-level handwritten text recognition

### v1

@@ III-B Document Attention Network
For each iteration $t$, the previously predicted tokens are put together as query input for the decoder: $\hat{\boldsymbol{y}}^{t}=[\hat{\boldsymbol{y}}_{0},\dots,\hat{\boldsymbol{y}}_{t-1}]$. A trainable weight matrix is used to embed each text token as a vector of dimension $C_{f}$, leading to $\boldsymbol{e}\in\mathbb{R}^{t\times C_{f}}$. Following the original transformer [28], one-dimensional positional encoding ($\boldsymbol{p}^{t}_{\text{1D}}$) is added, leading to the query sequence $\boldsymbol{q}^{t}=\boldsymbol{e}^{t}+\boldsymbol{p}^{t}_{\text{1D}}$. In the same way, its 2D adaptation [30] encoding both dimensions of the image on half of the channels is used to inject positional information to the image features, resulting in $\boldsymbol{g}=\boldsymbol{f}+\boldsymbol{p}_{\text{2D}}$. Query tokens $\boldsymbol{q}^{t}$ and images features $\boldsymbol{g}$ are processed through 8 stacked transformer decoder layers resulting in $\boldsymbol{o}^{t}=d(\boldsymbol{q}^{t},\boldsymbol{g})$, preserving the number of channels $C_{f}$. Finally, a score $\boldsymbol{s}_{t}$ for each token of the token set $\mathcal{A}$ is computed through a linear projection layer $p$ based on the output corresponding to the last token of the query sequence:
$\boldsymbol{s}_{t}=p(\boldsymbol{o}_{t-1}).$ (1)
The predicted token corresponds to the one with the highest score: $\hat{\boldsymbol{y}}_{t}=\operatorname*{arg\,max}(\boldsymbol{s}_{t})$.
$\mathcal{L}_{\text{DAN}}=\sum_{t=1}^{L+1}\mathcal{L}_{\text{CE}}(\boldsymbol{y}_{t},\boldsymbol{p}_{t}),$ (2)
with $\boldsymbol{p}_{t,i}=\operatorname*{soft\,max}(\boldsymbol{s}_{t})=\frac{e^{\boldsymbol{s}_{t,i}}}{\sum_{j}e^{\boldsymbol{s}_{t,j}}}$.
@@
[width=]images/tikz/dan_in_out DAN [13]
@@
[width=]images/tikz/fdan_in_out Faster DAN [14]
@@
[width=]images/tikz/wdan_in_out W-DAN (ours, $w=3$)
@@
[width=]images/tikz/mt-dan_in_out MT-DAN (ours, $m=3$)
@@
[width=0.928]images/tikz/meta_dan_in_out Meta-DAN (ours, $w=3$, $m=2$)
[…]
III-C Faster DAN
The Faster DAN [13] was proposed as an alternative to speed up the decoding process at inference stage. It has to be noted that the Faster DAN architecture is exactly the same as the DAN. The contribution is about decoding strategy whose aim is to parallelize the text line predictions. It relies on a two-stage approach. First, an autoregressive process predicts the first character of each line, following the original DAN strategy. Then, all lines are completed in parallel, reducing the total number of iterations. The main difference appears in that second decoding stage. Instead of using the output corresponding to the last token $\boldsymbol{o}_{t-1}$ only, the token prediction is performed using the outputs corresponding to the last token of each text line. In this way, for $N$ text lines, there are $N$ token predictions per iteration. Equation 1 becomes:
$\boldsymbol{s}^{l}_{t}=p(\boldsymbol{o}^{l}_{t-1}),$ (3)
where $\boldsymbol{s}^{l}_{t}$ is the score vector associated to the token prediction of the next character for line $l$ at iteration $t$, and $\boldsymbol{o}^{l}_{t-1}$ is the transformer output corresponding to the last predicted token for line $l$.
The Faster DAN introduces a specific positional encoding to associate each text token with a line index and a positioning index within that line. This is a crucial component to make this parallelization possible. Although the Faster DAN showed interesting inference speed improvements, it results in a loss of past context when predicting a character. Indeed, instead of benefiting from the whole past context as is the case for the DAN, the Faster DAN has a partial view of the past and future due to the parallel line predictions, which reduces its ability to model the language.
In this paper, we propose a strategy to get the best of both worlds: a reduced prediction time through query parallelization, and an improved context modeling by processing successive queries altogether and relying on multi-token predictions.
IV Meta Document Attention Network
We propose two new approaches for the autoregressive prediction of text tokens in the context of end-to-end handwritten text recognition at page level. For both, the idea is to predict multiple tokens at each decoding iteration to reduce the number of decoding steps, while preserving the whole past context. We assume this is a key element to reduce the number of recognition errors through language modeling. First, we propose to associate multiple predictions to a single query $q_{t-1}$, to predict the $m$ next tokens at once. This is referred as the Document Attention Network with Multi-Tokens predictions (MT-DAN). Second, we propose a query approach based on a window mechanism to process multiple queries at once, each associated with a single prediction. This is referred as the Document Attention Network with Windowed queries (W-DAN). Finally, we design a global strategy including these two approaches, called Meta Document Attention Network (Meta-DAN).
A visual comparison of the prediction processes is depicted in Figure III-B.
IV-A MT-DAN
@@
$\boldsymbol{s}^{(\boldsymbol{t})}_{t+k}=p^{k}(\boldsymbol{o}_{t-1}),$ (4)
where $\boldsymbol{s}^{(\boldsymbol{t})}_{t+k}$ is the score vector associated to the token prediction of the $k^{\text{th}}$ next token computed at iteration $t$, $p^{k}$ is the $k^{\text{th}}$ projection head, and $\boldsymbol{o}_{t-1}$ is the transformer output corresponding to the last query token $\boldsymbol{q}_{t-1}$. In this way, $m$ predictions are performed per iteration following:
@@
As one can note, several predictions $\hat{\boldsymbol{y}}^{(\boldsymbol{t})}_{t+k}$ will be associated to the same target token $\boldsymbol{y}_{t+k}$. This is not an issue during training, we train the MT-DAN using the cross-entropy loss over the whole token sequence, as for the original DAN. It involves adding $m-1$ extra <e> tokens in order for all text tokens to have a target for each of the projection heads, leading to the following loss:
$\mathcal{L}_{\text{MT-DAN{}}}=\sum_{t=1}^{L+1}\sum_{k=0}^{m-1}\mathcal{L}_{\text{CE}}(\boldsymbol{y}_{t+k},\boldsymbol{p}^{(\boldsymbol{t})}_{t+k}),$ (6)
@@
In Section VI, we propose a decoding policy based on the confidence level of the prediction to choose not to predict all $m$ next tokens for poor confidence. We demonstrate it is a good trade-off between prediction speed and recognition performance.
Predicting multiple tokens $\hat{\boldsymbol{y}}_{t+k}$ from a unique output $\boldsymbol{o}_{t-1}$ can theoretically lead to saturation as $m$ increases. Indeed, the fixed-size vector $\boldsymbol{o}_{t-1}$ must represent the next $m$ text tokens at once. To solve this issue, we propose another way to predict multiple tokens per iteration by processing several consecutive queries at once.
IV-B W-DAN
[…]
[width=]images/tikz/dan_attention DAN [13]/MT-DAN (ours)
[…]
[width=]images/tikz/fdan_attention Faster DAN [14]
[…]
[width=]images/tikz/wdan_attention W-DAN (ours) / Meta-DAN (ours)
(for $w=3$)
[…]
To solve the MT-DAN output compression issue, we propose the Document Attention Network with Windowed queries (W-DAN). Instead of relying on a single output to predict $m$ tokens, we opt for using the output of the last $w$ predictions to predict the $w$ next tokens, $w$ being a new hyperparameter. In this way, it solves the compression issue as they are as many output as they are predictions. More precisely, each output $\boldsymbol{o}_{t-q}$ is responsible for the prediction of token $\hat{\boldsymbol{y}}_{t-q+w}$, with $q\in[1,\dots,w]$. Equation 1 becomes:
$\boldsymbol{s}_{t-q+w}=p(\boldsymbol{o}_{t-q}),$ (7)
@@ IV-B W-DAN
[…]
$\mathcal{L}_{\text{W-DAN{}}}=\sum_{t=w}^{L+1+n_{\text{e}}}\mathcal{L}_{\text{CE}}(\boldsymbol{y}_{t},\boldsymbol{p}_{t}),$ (9)
There is a major advantage of both presented approaches compared to the DAN and the Faster DAN: the $k$ or $w$ predictions follow one another and are performed altogether. It means that the context is enlarged when prediction takes place. The model must focus on a broader context to predict multiple successive tokens. For instance, for $k=3$ (or $w=3$), the model focuses on predicting $\{\hat{\boldsymbol{y}}_{t},\hat{\boldsymbol{y}}_{t+1},\hat{\boldsymbol{y}}_{t+2}\}$ at once. This way, the prediction of $\hat{\boldsymbol{y}}_{t}$ is impacted by the future context $\{\hat{\boldsymbol{y}}_{t+1},\hat{\boldsymbol{y}}_{t+2}\}$, which improves the ability to model the language at prediction time compared with DAN (which only benefits from past context). The same goes with $\hat{\boldsymbol{y}}_{t+1}$ predicted with $\{\hat{\boldsymbol{y}}_{t+2}\}$ as context of future.
IV-C Meta-DAN
We propose now to formulate a generic prediction strategy to take advantage of both presented approaches by combining them. It is referred as the Meta Document Attention Network (Meta-DAN). Thus, the Meta-DAN is parametrized by two hyperparameters $m$ and $w$, indicating the number of projection heads and the window size, respectively. In other words, at each iteration $t$, a window of $w$ query tokens $\{\boldsymbol{q}_{t-w},\dots,\boldsymbol{q}_{t-1}\}$ are processed at once, each one leading to $m$ predictions. More precisely, the output of the query token $\boldsymbol{q}_{t-q}$, i.e. $\boldsymbol{o}_{t-q}$, is responsible for the prediction of $\{\hat{\boldsymbol{y}}_{t-q+w},\dots,\hat{\boldsymbol{y}}_{t-q+w+m-1}\}$ as follows:
$\hat{\boldsymbol{y}}^{(\boldsymbol{t})}_{t}=\operatorname*{arg\,max}(\boldsymbol{s}^{(\boldsymbol{t})}_{t}),$ (10)
@@ IV-C Meta-DAN
$\boldsymbol{s}^{(\boldsymbol{t-q+1})}_{t-q+w+k}=p^{k}(\boldsymbol{o}_{t-q}),$ (11)
where $\boldsymbol{s}^{t-q+1}_{t-q+w+k}$ is the score associated to token $\hat{\boldsymbol{y}}_{t-q+w+k}$ by the $k^{\text{th}}$ projection head of the $q^{\text{th}}$ query token in the query window, and with $k\in[0,\dots,m-1]$ and $q\in[1,\dots,w]$. The Meta-DAN involves initializing the decoding process with $w$ <s> tokens, as for the W-DAN, and adding extra <e> so that the last query has a target prediction for all its associated projection heads.
Regarding training, equation 2 becomes:
$\mathcal{L}_{\text{Meta-DAN{}}}=\sum_{t=w}^{L+1+n_{\text{e}}}\sum_{k=0}^{m-1}\mathcal{L}_{\text{CE}}(\boldsymbol{y}_{t+k},\boldsymbol{p^{(\boldsymbol{t})}_{t+k}})$ (12)
Detailed features of evaluated datasets (page-level). Dataset Train Val Test Charset Average Language Century Writers size # char. per page IAM (paragraph) [37] 747 161 336 80 371 English 20th-21st 657 RIMES-2009 [38] 1,050 100 100 108 588 French 21st $\approx 400$ RIMES-2024 [39] 4,500 505 600 123 567 French 21st $\approx$ 1,400 READ-2016 [40] 350 50 50 89 468 German 15th-19th unknown Esposalles [41] 75 25 25 62 1,748 Catalan 17th 1 BRESSAY [42] 647 154 199 109 2,612 Portuguese 21st 1,000 ScribbleLens [43] 164 14 21 85 1,410 Dutch 16th-18th 85 Eparchos [44] 80 20 20 157 962 Greek 16th 2 CASIA-2 [45] 3,600 475 1,015 2,704 273 Chinese 21st 1,019 MAURDOR-C3 [46] 1,006 148 166 134 491 French/English 21st > 130 MAURDOR-C4 [46] 721 111 114 127 675 French/English 21st > 90
As one can note, several query tokens can be in charge of the same prediction. This is shown in Figure III-B with a Meta-DAN example with $w=3$ and $m=2$: the "m" and "p" are predicted twice, from two different queries. In this work, we limit ourselves to use the prediction of the first projection head for the $w-1$ first queries, and the prediction of all projection heads for the last query. While discarded, we assume that training to predict those duplicated characters encourages the network to use its future context to improve language modeling. In this way, all queries follow the same behavior, which make training easier.
In summary, the Meta-DAN approach is a general formulation for which one can distinguish three specific cases depending on the hyperparameters:
@@
IV-D Impact on the computational complexity
@@
To further give insights about the main differences between all these prediction strategies, Figure IV-B provide visuals of the attending maps (transformer masks), stating which token can access to which other tokens.
@@
Regarding the MT-DAN approach, the attending maps of the DAN and the MT-DAN are identical; this means that they both imply the same computations in the transformer decoder layers. In addition, contrary to the NLP application which relies on the Byte-Pair Encoding algorithm leading to large vocabularies, the addition of extra projection heads do not significantly impact the global computational complexity as we are working at character level, limiting the dimensions of the projection weight matrices.
@@
Concerning the W-DAN approach (and thus the Meta-DAN), the attending maps show that there are more interactions between the tokens in the self-attention layers, resulting in more computations, theoretically. However, to our knowledge, the current implementations of the transformer relies on masks to "hide" the forbidden tokens (usually the future) after the query-key multiplication takes place. In this way, this does not impact the computation time in practice since computations are performed over all token pairs whether they are really used or not.
@@
No matter the approach (DAN, Faster DAN, MT-DAN, W-DAN, Meta-DAN), the prediction process can be parallelized at training time using teacher forcing, so there is no additional benefit for the training stage. As stated previously, the number of computations are not significantly impacted neither by the MT-DAN nor by the W-DAN. The major impact is positive and related to the time complexity at prediction time when parallelizing the computation of token predictions.
V Experimental conditions
V-A Datasets
@@
We evaluate the proposed approach on many public HTR datasets, namely IAM [37], READ-2016 [40], Esposalles [41], BRESSAY [42], ScribbleLens [43], Esparchos [44], CASIA-2 [45], MAURDOR [46], RIMES-2009 [38] and RIMES-2024 [39]. They cover 8 languages, three alphabets (Latin, Greek and Chinese), 6 centuries (15th-21st). It enables us to have a robust evaluation due to the high diversity of the datasets. Indeed, as detailed in Table IV-C, they involve many writing configurations (from single writer to more than a thousand), various sizes of character set, as well as a variable amount of training data (from less than a hundred to few thousand pages).
@@
We used the standard split into training, validation and test sets when existing. To our knowledge, we are the first to use the recently released RIMES-2024 dataset, so we propose a new split, randomly selecting samples for training, validation and test by respecting the split provided at line level in 2011 [47]. We discarded images from the "Q" category, which corresponds to forms, since they do not provide the whole transcript. For multilingual experiments, we discard the RIMES-2009 dataset from the evaluation setting as some test samples can be found in the RIMES-2024 training set. For ScribbleLens, we removed 10% of the training samples to create a validation set by randomly picking from each of its data sources. For Eparchos, we propose a split at page-level by taking 80% of the samples for training, 10% for validation and 10% for test. Appendix B provides the splits at line level, used for pre-training purposes, for the datasets providing line-level annotations. This section also includes some input document image examples.
[b] Comparison of W-DAN recognition performance with respect to the window size $w$ on the IAM and BRESSAY test sets. Speed Factor (SF) is indicated as a comparison with the DAN baseline.
Window IAM BRESSAY
size ($w$) CER (%) WER (%) Time (s) SF CER (%) WER (%) Time (s) SF
1 (DAN) 4.23 14.00 3.39 x1.0 2.43 7.85 24.65 x1.0
2 4.67 15.14 1.94 x1.7 2.91 8.87 20.86 x1.2
3 4.83 15.62 1.26 x2.7 3.05 9.24 12.25 x2.0
4 4.75 15.71 0.97 x3.5 2.63 9.35 8.63 x2.8
5 4.87 16.28 0.77 x4.4 3.02 9.72 6.63 x3.7
10* 5.07 16.53 0.40 x8.5 2.66 9.53 2.95 x8.4
25* 6.56 20.94 0.18 x18.8 4.37 15.21 1.13 x21.8
    Training was performed over 2M and 3M training samples for $w=10$ and $w=25$, respectively (instead of 1M).
V-B Training strategy

### v2

@@ 4 Meta Document Attention Network
[width=]dan_in_out DAN [14]
@@
[width=]fdan_in_out Faster DAN [16]
@@
[width=]wdan_in_out W-DAN ($w=3$)
@@
[width=]mt-dan_in_out MT-DAN ($m=3$)
@@
[width=0.85]meta_dan_in_out Meta-DAN ($w=3$, $m=2$)
@@
4.1 MT-DAN
@@
$\boldsymbol{s}^{(\boldsymbol{t-1})}_{t+k}=\boldsymbol{W}^{k}\boldsymbol{o}_{t-1},$ (3)
where $\boldsymbol{s}^{(\boldsymbol{t-1})}_{t+k}$ is the score vector associated with the token prediction of the $k^{\text{th}}$ next token computed at iteration $t$. The exponent $(\boldsymbol{t-1})$ indicates that this prediction was performed based on query indexed by t1t-1. $\boldsymbol{W}^{k}$ is the weight matrix corresponding to the $k^{\text{th}}$ projection head, and $\boldsymbol{o}_{t-1}$ is the transformer output corresponding to the last query token $\boldsymbol{q}_{t-1}$. In this way, $m$ predictions are performed per iteration following:
@@
As one can note, several predictions $\hat{\boldsymbol{y}}^{(\boldsymbol{t})}_{t+k}$ are associated with the same target token $\boldsymbol{y}_{t+k}$. This is not an issue during training; we train the MT-DAN using the cross-entropy loss over the whole token sequence, as for the original DAN. It involves adding m1m-1 extra <e> tokens in order for all text tokens to have a target for each of the projection heads, leading to the following loss:
$\mathcal{L}_{\text{MT-DAN{}}}=\sum_{t=1}^{L^{y}+1}\sum_{k=0}^{m-1}\mathcal{L}_{\text{CE}}(\boldsymbol{y}_{t+k},\boldsymbol{p}^{(\boldsymbol{t})}_{t+k})$, (5)
@@
Predicting multiple tokens $\hat{\boldsymbol{y}}_{t+k}$ from a unique output $\boldsymbol{o}_{t-1}$ can theoretically lead to saturation as $m$ increases. Indeed, the fixed-size vector $\boldsymbol{o}_{t-1}$ must represent the next $m$ text tokens at once. To solve this issue, we propose another way to predict multiple tokens per iteration, which consists in processing several consecutive queries at once.
4.2 W-DAN
To solve the MT-DAN output compression issue, we propose the Document Attention Network with Windowed queries (W-DAN). Instead of relying on a single output to predict $m$ tokens, we opt for using the output of the last $w$ predictions to predict the $w$ next tokens, $w$ being a new hyperparameter. In this way, it solves the compression issue as there are as many outputs as there are predictions. The decoding process of the W-DAN acts as a sliding window over the query tokens, processing them by chunks of $w$ tokens at once, leading to $w$ new outputs. More precisely, each output $\boldsymbol{o}_{t-q}$ is responsible for the prediction of a single token $\hat{\boldsymbol{y}}_{t-q+w}$, with $q\in[1,\dots,w]$. Equation 1 becomes:
$\boldsymbol{s}_{t-q+w}=\boldsymbol{W}\boldsymbol{o}_{t-q},$ (6)
@@ 4.2 W-DAN
It has to be noted that all scores $\boldsymbol{s}_{t-q+w}$ are computed using the same projection head of weights $\boldsymbol{W}$ as they are applied independently on each $\boldsymbol{o}_{t-q}$. In addition, it is necessary to initialize the decoding process with $w$ initial tokens to predict the first $w$ tokens. We arbitrarily choose to use the same start-of-transcription token $w$ times ($\hat{\boldsymbol{y}}_{0}=\cdots=\hat{\boldsymbol{y}}_{w-1}=\text{<s>}$). We also added $n_{\text{e}}$ extra end-of-transcription tokens at the end of the target sequence so as to make the target length divisible by $w$. It leads to the following loss:
$\mathcal{L}_{\text{W-DAN{}}}=\sum_{t=w}^{L^{y}+1+n_{\text{e}}}\mathcal{L}_{\text{CE}}(\boldsymbol{y}_{t},\boldsymbol{p}_{t}).$ (8)
The major advantage of both presented approaches compared to the DAN and the Faster DAN is that the $k$ or $w$ predictions are consecutive and performed altogether: the context is enlarged when the prediction takes place. The model must focus on a broader context to predict multiple successive tokens. For instance, for $k=3$ (or $w=3$), the model focuses on predicting $\{\hat{\boldsymbol{y}}_{t},\hat{\boldsymbol{y}}_{t+1},\hat{\boldsymbol{y}}_{t+2}\}$ at once. This way, the prediction of $\hat{\boldsymbol{y}}_{t}$ is impacted by the future context $\{\hat{\boldsymbol{y}}_{t+1},\hat{\boldsymbol{y}}_{t+2}\}$, which improves the ability to model the language at prediction time compared to DAN.
4.3 Meta-DAN
We propose to take advantage of both presented approaches by combining them: this is referred to as the Meta Document Attention Network (Meta-DAN). Thus, the Meta-DAN is parametrized by two hyperparameters $m$ and $w$, indicating the number of projection heads and the window size, respectively. In other words, at each iteration $t$, a window of $w$ query tokens $\{\boldsymbol{q}_{t-w},\dots,\boldsymbol{q}_{t-1}\}$ are processed at once, each one leading to $m$ predictions. More precisely, the output of the query token $\boldsymbol{q}_{t-q}$, i.e. $\boldsymbol{o}_{t-q}$, is responsible for the prediction of $\{\hat{\boldsymbol{y}}_{t-q+w},\dots,\hat{\boldsymbol{y}}_{t-q+w+m-1}\}$ as follows:
$\hat{\boldsymbol{y}}^{(\boldsymbol{t-q})}_{t-q+w+k}=\operatorname*{arg\,max}(\boldsymbol{s}^{(\boldsymbol{t-q})}_{t-q+w+k})$, (9)
@@ 4.3 Meta-DAN
$\boldsymbol{s}^{(\boldsymbol{t-q})}_{t-q+w+k}=\boldsymbol{W}^{k}\boldsymbol{o}_{t-q},$ (10)
where $\boldsymbol{s}^{(\boldsymbol{t-q})}_{t-q+w+k}$ is the score associated with token $\hat{\boldsymbol{y}}_{t-q+w+k}$ by the $k^{\text{th}}$ projection head of the $q^{\text{th}}$ query token in the query window, and with $k\in[0,\dots,m-1]$ and $q\in[1,\dots,w]$. The Meta-DAN involves initializing the decoding process with $w$ <s> tokens, as for the W-DAN, and adding $n_{e}$ extra <e> so that the last query has a target prediction for all its associated projection heads. Regarding training, equation 2 becomes:
$\mathcal{L}_{\text{Meta-DAN{}}}=\sum_{t=w}^{L^{y}+1+n_{\text{e}}}\sum_{k=0}^{m-1}\mathcal{L}_{\text{CE}}(\boldsymbol{y}_{t+k},\boldsymbol{p}^{(\boldsymbol{t})}_{t+k})$ (11)
As one can note, several query tokens can be in charge of the same prediction. This is shown in Figure 4 with an example with $w=3$ and $m=2$: the "m" and "p" are predicted twice, from two different queries. In this work, we limit ourselves to using the prediction of the first projection head for the w1w-1 first queries, and the prediction of all projection heads for the last one. Although discarded, we assume that training to predict these extra characters encourages the network to improve language modeling.
In summary, the Meta-DAN is a general formulation for which one can distinguish three specific cases depending on the hyperparameters:
@@
4.4 Impact on the computational complexity
@@
To further give insights into the main differences between all these prediction strategies, Figure 4.4 provide visuals of the attending maps (transformer masks), stating which token can access to which other tokens for the masked self-attention.
[…]
[width=]dan_attention DAN [14]/MT-DAN (ours)
[…]
[width=]fdan_attention Faster DAN [16]
[…]
[width=]wdan_attention W-DAN / Meta-DAN
(ours, for $w=3$)
[…]
@@
Regarding the MT-DAN approach, the attending maps of the DAN and the MT-DAN are identical, no matter the value of $m$; this means that they both imply the same computations in the transformer decoder layers. In addition, contrary to the NLP application, which relies on the Byte-Pair Encoding algorithm leading to large vocabularies, the addition of extra projection heads does not significantly impact the global computational complexity as we are working at character level, limiting the dimensions of the projection weight matrices.
@@
Concerning the W-DAN approach (and thus the Meta-DAN), the attending maps show that there are more interactions between the tokens in the self-attention layers, resulting in more computations, theoretically. However, to our knowledge, the current implementations of the transformer rely on masks to "hide" the forbidden tokens (usually the future) after the query-key multiplication takes place. In this way, this does not impact the computation time in practice since computations are performed over all token pairs, whether they are really used or not.
@@
No matter the approach, the prediction process can be parallelized at training time using teacher forcing [46], i.e., by injecting the whole ground truth instead of the autoregressive previous predictions, so that all decoding iterations are performed at once. In this way, there is no extra benefit for the training stage. The major impact is positive and related to the time complexity at prediction time when parallelizing the computation of token predictions.
5 Experimental conditions
5.1 Datasets
@@
We select 10 public datasets with high diversity to make our evaluation robust: IAM [31], READ-2016 [35], Esposalles [34], BRESSAY [17], ScribbleLens [19], Esparchos [42], CASIA-2 [30], MAURDOR [6], RIMES-2009 [27] and RIMES-2024 [28]. As detailed in Table 5.1, they cover 8 languages, three alphabets (Latin, Greek, and Chinese), and 6 centuries. They also involve many writing configurations (from a single writer to more than a thousand), various sizes of character sets, as well as a variable amount of training data.
Detailed features of evaluated datasets (page-level). Dataset Train Val Test Charset Average Language Century Writers size # char/page IAM (paragraph) [31] 747 161 336 80 371 English 20th-21st 657 RIMES-2009 [27] 1,050 100 100 108 588 French 21st $\approx 400$ RIMES-2024 [28] 4,500 505 600 123 567 French 21st $\approx$ 1,400 READ-2016 [35] 350 50 50 89 468 German 15th-19th unknown Esposalles [34] 75 25 25 62 1,748 Catalan 17th 1 BRESSAY [17] 647 154 199 109 2,612 Portuguese 21st 1,000 ScribbleLens [19] 164 14 21 85 1,410 Dutch 16th-18th 85 Eparchos [42] 80 20 20 157 962 Greek 16th 2 CASIA-2 [30] 3,600 475 1,015 2,704 273 Chinese 21st 1,019 MAURDOR-C3 [6] 1,006 148 166 134 491 French/English 21st > 130 MAURDOR-C4 [6] 721 111 114 127 675 French/English 21st > 90
@@
We used the standard split into training, validation, and test sets when existing. To our knowledge, we are the first to use the recently released RIMES-2024 dataset, so we propose a split, randomly sampling for training, validation and test by respecting the split provided at line level in 2011 [26]. We discarded images from the "Q" category, which corresponds to forms, since they do not provide the whole transcript. For ScribbleLens, we removed 10% of the training samples to create a validation set by randomly picking from each of its data sources. For Eparchos, we propose a split at page level, by taking 80% of the samples for training, 10% for validation, and 10% for test. A provides the splits at line level, used for pre-training purposes, for the datasets providing line-level annotations. This section also includes document image examples.
5.2 Training strategy